OpenStudy (anonymous):

How could I solve 3x-5y+7, 5x-2y=-1 This is the one question I just cannot get right.

OpenStudy (anonymous):

Its a system of equations

OpenStudy (anonymous):

You solve for a variable in one of them, and then plug that into the other for whatever variable you solve for

OpenStudy (anonymous):

Need me to get you started?

OpenStudy (anonymous):

I know how to start it thanks, but I always end up getting some weird result. If you could quickly show the steps that would be great thanks.

OpenStudy (anonymous):

So okay, does that mean I can just pick either of the equations and solve for either x or y?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

The first equation doesn't have an equal sign. Is it a typo? Can you correct please?

OpenStudy (anonymous):

Sorry, 3x-5y=7

OpenStudy (anonymous):

Ah, okay then let's say I pick the second equation to solve first. Then 5x-2y=-1 x = (2y-1)/5 do you agree?

OpenStudy (anonymous):

Yes, I've got that thanks

OpenStudy (anonymous):

Then I would use X = to substitute into first eq. (3/5)(2y-1) -5y =7 Personally, I don't work well with fractions so I would multiply all terms by (5/1)

OpenStudy (anonymous):

Then i get 3(2y-) -25y =35

OpenStudy (anonymous):

oops typo! 3(2y-1)-25y = 35

OpenStudy (anonymous):

Now 6y -3 -25y = 35 Gather like terms solve for y. -19y = 38 Does that look right?

OpenStudy (anonymous):

Thank you, I see where I kept going wrong. Thanks!

OpenStudy (anonymous):

No prob-glad to help!